Dorchester Company, on March 1, 2019 has a beginning Work in Process inventory of zero. All materials are added into production at the beginning of its production. There is only one production WIP inventory. On March 1, Dorchester started into production 18,500 units. At the end of the month there were 13,000 units completed and transferred into the Finished Goods Inventory. The ending WIP was 40% complete with respect to conversion. For the month of March the following costs were incurred and recorded in the WIP: Direct Material $11,000 Direct Labor 27,000 Factory Overhead 25,000 Dorchester uses the weighted-average process costing method. Use this information to determine the cost per equivalent unit of conversion for the month of March: (Round & enter final answers to the nearest cent.)
Answer #1

Equivalent unit of conversion = Unit transferred out+(Ending WIP*percent completion)

= 13000+(5500*40%)

Equivalent unit of conversion = 15200

Cost per equivalent unit of conversion = (27000+25000)/15200 = 3.42